OpenStudy (anonymous):

NEED HELP WITH EXPONENTIAL AND LOGARITHMIC FUNCTIONS log(2x+3)=3 help me understand??

OpenStudy (anonymous):

you are supposed to solve for \(x\) right?

OpenStudy (anonymous):

yes i believe so.

OpenStudy (anonymous):

assuming this is log base ten, then \[\log(x)=y\iff 10^y=x\]

OpenStudy (anonymous):

why would we assume it is 10?? do we always do that?

OpenStudy (anonymous):

they are just different ways of saying the same thing so if \[\log(2x+3)=3\] that means \[2x+3=10^3\] and you can solve for \(x\)

OpenStudy (anonymous):

ya give me a secccc

OpenStudy (anonymous):

if you don't see a base written, like \(\log(x)\) instead of \(\log_2(x)\) then you assume it is log base ten

OpenStudy (anonymous):

x=498.5?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

\[2x+3=1000\\ 2x=997\\ x=997\div 2\] whatever that is

OpenStudy (anonymous):

yea thats exatly what my work looks like!!

OpenStudy (anonymous):

one of the answer choices is 997/2 so im guessing they dont want us to simplify??

OpenStudy (anonymous):

k that was the routine part the main part was to be able to go from logarithmic to exponential form quickly which is what you will need \[\large b^y=x\iff \log_b(x)=y\]

OpenStudy (anonymous):

yeah i would go with \(\frac{997}{2}\)

OpenStudy (anonymous):

awesome! You are definitely going to go far in life! You are so smart and helpful!! thank you so much :)

OpenStudy (anonymous):

lol thanks and your welcome